OpenStudy (anonymous):

Find the average rate of change of f(x)=1+sin x over interval [0,pi]

OpenStudy (anonymous):

Average rate of change of a function \(f\) over some interval \([a,b]\), provided that \(f\) is continuous over the interval, is \[\frac{f(b)-f(a)}{b-a}\]

OpenStudy (anonymous):

you got this?

OpenStudy (isaiah.feynman):

@SithsAndGiggles Same thing came to my head but what is a? it is 0?

OpenStudy (anonymous):

Yes, \([a,b]=[0,\pi]\), so \(a=0\) and \(b=\pi\).

OpenStudy (isaiah.feynman):

OK. Now I'm calm. Oh you forgot to add the limit as b approaches a.

OpenStudy (anonymous):

That's the case for instantaneous rate of change. We just want the average rate of change.

OpenStudy (isaiah.feynman):

Oh lol I didn't read the question well.